# BrainSCUBA: Fine-Grained Natural Language Captions of Visual Cortex Selectivity

### v1

@@ Abstract
Understanding the functional organization of higher visual cortex is a central focus in neuroscience. Past studies have primarily mapped the visual and semantic selectivity of neural populations using hand-selected stimuli, which may potentially bias results towards pre-existing hypotheses of visual cortex functionality. Moving beyond conventional approaches, we introduce a data-driven method that generates natural language descriptions for images predicted to maximally activate individual voxels of interest. Our method – Semantic Captioning Using Brain Alignments (“BrainSCUBA”) – builds upon the rich embedding space learned by a contrastive vision-language model and utilizes a pre-trained large language model to generate interpretable captions. We validate our method through fine-grained voxel-level captioning across higher-order visual regions. We further perform text-conditioned image synthesis with the captions, and show that our images are semantically coherent and yield high predicted activations. Finally, to demonstrate how our method enables scientific discovery, we perform exploratory investigations on the distribution of “person” representations in the brain, and discover fine-grained semantic selectivity in body-selective areas. Unlike earlier studies that decode text, our method derives voxel-wise captions of semantic selectivity. Our results show that BrainSCUBA is a promising means for understanding functional preferences in the brain, and provides motivation for further hypothesis-driven investigation of visual cortex.
@@ 2 Related Work
Several recent studies have yielded intriguing results by using large-scale vision-language models to reconstruct images and text-descriptions from brain patterns when viewing images (Takagi & Nishimoto, 2022; Chen et al., 2022; Doerig et al., 2022; Ferrante et al., 2023; Ozcelik & VanRullen, 2023; Liu et al., 2023), or to generate novel images that are predicted to activate a given region (Ratan Murty et al., 2021; Gu et al., 2022; Luo et al., 2023). Broadly speaking, these approaches require conditioning on broad regions of the visual cortex, and have not demonstrated the ability to scale down and enable voxel-level understanding of neural selectivity. Additionally, these methods produce images rather than interpretable captions. Below we review some of this work in more detail.
@@
There are two broad directions when it comes to brain conditioned generative models for vision. The first seeks to decode (reconstruct) visual inputs from the corresponding brain acitvations, including works that leverage retrieval, variational autoencoders (VAEs), generative adversarial networks (GANs), and score/energy/diffusion models (Kamitani & Tong, 2005; Han et al., 2019; Seeliger et al., 2018; Shen et al., 2019; Ren et al., 2021; Takagi & Nishimoto, 2022; Chen et al., 2023; Lu et al., 2023; Ozcelik & VanRullen, 2023). Some approaches further utilize or generate captions that describe the observed visual stimuli (Doerig et al., 2022; Ferrante et al., 2023; Liu et al., 2023; Mai & Zhang, 2023).
@@ 4 Results
In this section, we utilize BrainSCUBA to generate voxel-wise captions and demonstrate that it can capture the selectivity in different semantic regions in the brain. We first show that the generated nouns are interpretable across the entire brain and exhibit a high degree of specificity within pre-identified category-selective regions. Subsequently, we use the captions as input to text-to-image diffusion models to generate novel images, and confirm the images are semantically consistent within their respective regions. Finally, we utilize BrainSCUBA to analyze the distribution of person representations across the brain. These results illustrate BrainSCUBA’s ability to characterize human visual cortical populations, rendering it a promising framework for exploratory neuroscience.
@@ 4.3 Text-Guided Brain Image Synthesis
Visualization of the captions can be helpful in highlighting subtle co-occurrence statistics, thereby enriching our understanding of the captions corresponding to images that activate a given region. We utilize a text-to-image diffusion model, and condition the synthesis process on the voxel-wise captions within an ROI (Figure 5).
@@ A.6 Human study details
For each question, a human evaluator would perform comparisons, from splits and the NSD subjects; with human evaluators per question. For the human study results in section 4.4, we collected total responses for the four questions.

### v3

@@ Abstract
Understanding the functional organization of higher visual cortex is a central focus in neuroscience. Past studies have primarily mapped the visual and semantic selectivity of neural populations using hand-selected stimuli, which may potentially bias results towards pre-existing hypotheses of visual cortex functionality. Moving beyond conventional approaches, we introduce a data-driven method that generates natural language descriptions for images predicted to maximally activate individual voxels of interest. Our method – Semantic Captioning Using Brain Alignments (“BrainSCUBA”) – builds upon the rich embedding space learned by a contrastive vision-language model and utilizes a pre-trained large language model to generate interpretable captions. We validate our method through fine-grained voxel-level captioning across higher-order visual regions. We further perform text-conditioned image synthesis with the captions, and show that our images are semantically coherent and yield high predicted activations. Finally, to demonstrate how our method enables scientific discovery, we perform exploratory investigations on the distribution of “person” representations in the brain, and discover fine-grained semantic selectivity in body-selective areas. Unlike earlier studies that decode text, our method derives voxel-wise captions of semantic selectivity. Our results show that BrainSCUBA is a promising means for understanding functional preferences in the brain, and provides motivation for further hypothesis-driven investigation of visual cortex.https://www.cs.cmu.edu/~afluo/BrainSCUBA
@@ 2 Related Work
Several recent studies have yielded intriguing results by using large-scale vision-language models to reconstruct images and text-descriptions from brain patterns when viewing images (Takagi & Nishimoto, 2022; Chen et al., 2022; Doerig et al., 2022; Ferrante et al., 2023; Ozcelik & VanRullen, 2023; Liu et al., 2023), or to generate novel images that are predicted to activate a given region (Ratan Murty et al., 2021; Gu et al., 2022; Luo et al., 2023). Broadly speaking, these approaches require conditioning on broad regions of the visual cortex, and have not demonstrated the ability to scale down and enable voxel-level understanding of neural selectivity. Additionally, these methods produce images rather than interpretable captions. Work on artificial neurons (Borowski et al., 2020; Zimmermann et al., 2021) have shown that feature visualization may not be more informative than top images in artificial neural networks. In contrast, our work tackles biological networks which have more noisy top-images that are less conducive to direct analysis, and the synthesis of novel images/captions can act as a source of stimuli for future hypothesis-driven neuroscience studies.
@@
There are two broad directions when it comes to brain conditioned generative models for vision. The first seeks to decode (reconstruct) visual inputs from the corresponding brain activations, including works that leverage retrieval, variational autoencoders (VAEs), generative adversarial networks (GANs), and score/energy/diffusion models (Kamitani & Tong, 2005; Han et al., 2019; Seeliger et al., 2018; Shen et al., 2019; Ren et al., 2021; Takagi & Nishimoto, 2022; Chen et al., 2023; Lu et al., 2023; Ozcelik & VanRullen, 2023). Some approaches further utilize or generate captions that describe the observed visual stimuli (Doerig et al., 2022; Ferrante et al., 2023; Liu et al., 2023; Mai & Zhang, 2023; Scotti et al., 2024).
@@ 4 Results
In this section, we utilize BrainSCUBA to generate voxel-wise captions and demonstrate that it can capture the selectivity in different semantic regions in the brain. We first show that the generated nouns are interpretable across the entire brain and exhibit a high degree of specificity within pre-identified category-selective regions. Subsequently, we use the captions as input to text-to-image diffusion models to generate novel images, and confirm the images are semantically consistent within their respective regions. Finally, we utilize BrainSCUBA to analyze the distribution of person representations across the brain to offer novel neuroscientific insight. These results illustrate BrainSCUBA’s ability to characterize human visual cortical populations, rendering it a promising framework for exploratory neuroscience.
@@ 4.3 Text-Guided Brain Image Synthesis
Visualization of the captions can be helpful in highlighting subtle co-occurrence statistics, with novel images critical for future hypothesis driven investigations of the visual cortex (Ratan Murty et al., 2021; Gu et al., 2023; Jain et al., 2023). We utilize a text-to-image diffusion model, and condition the synthesis process on the voxel-wise captions within an ROI (Figure 5).
@@ Conclusion.
6 Ethics
Institutional review board approval was requested and granted for the human study on body region images. The NSD dataset contains 8 subjects from age 19 to 32; half white, half asian; 2 men, 6 women. These demographics are not necessarily reflective of the broader population. As part of our framework, we use a pretrained language model based on fine-tuned GPT-2, and will reflect the biases present in the training and fine-tuning data. We believe language models trained on more diverse datasets, or those trained with weaker constraints will help mitigate issues of bias. We also note that text-conditioned image diffusion models can reflect negative stereotypes, but such issues could be mitigated on a case-by-case basis using concept editing (Gandikota et al., 2024).
7 Acknowledgments
This work used Bridges-2 at Pittsburgh Supercomputing Center through allocation SOC220017 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program, which is supported by National Science Foundation grants #2138259, #2138286, #2138307, #2137603, and #2138296. We also thank the Carnegie Mellon University Neuroscience Institute for support.
@@ Appendix A Appendix
8.
Top adjectives and more sentences (A.8)
9.
Encoder fitting stability (A.9)
10.
Ground truth functional localizer category distribution (A.10)
11.
Fine-grained concept distribution outside EBA (A.11)
12.
Norm of the embeddings with and without decoupled projection (A.12)
@@ A.6 Human study details
For the human study results in section 4.4, a human evaluator would perform comparisons, from splits and the NSD subjects; with human evaluators per question. We collected total responses for the four questions in the main text.
For the human study results below, a human evaluator would perform judgements, from splits and the NSD subjects; with human evaluators per question; across the clusters of images. We collected total responses for the eight questions below.
@@ A.7 Training and inference details
A.8 Top adjectives and more sentences
A.9 Encoder fitting stability
A.10 Ground truth functional localizer category distribution
A.11 Fine-grained concept distribution outside EBA
A.12 Norm of the embeddings with and without decoupled projection